# DC-VSR: Spatially and Temporally Consistent Video Super-Resolution with Video Diffusion Prior

### v1

@@ 4.2. Component Evaluation
Table 2. Ablation study of SAP & TAP and comparison with previous guidance methods using REDS4 dataset. $\text{DOVER}^{*}$ denotes DOVER ($\times$100). The best and second-best scores are written as bold and underline. Our proposed DSSAG is 1.5x faster than SAG and PAG.
@@
Figure 6. Qualitative comparison with previous guidance approaches.
@@ 5. Conclusion
Since DC-VSR is based on a diffusion model, it may generate false details and face difficulties in achieving real-time VSR. Moreover, DC-VSR may struggle to maintain temporal consistency across distant temporal tiles in long video clips, as TAP primarily relies on nearby tiles. Addressing this issue could be a promising future direction.
@@
We have attached a video named “SuppleVideo.mp4” as Supplementary Material, containing VSR results and comparisons that include the contents mentioned in the main paper.
@@ Appendix A Implementation Details
We fine-tune Image-to-Video Stable Video Diffusion (I2V-SVD) (StabilityAI, 2023) with 8×8$\times$8 × downsampling VAE (Rombach et al., 2022), trained on 14-frame sequences at a resolution of $576\times 1024$. We train DC-VSR using AdamW optimizer (Loshchilov and Hutter, 2019) with a learning rate of $1\times 10^{-4}$ and a batch size of 256 on 4 NVIDIA A100-80GB GPUs. Due to the memory limit, we use gradient accumulation to match the size of the batch. The training data is cropped into spatio-temporal tiles with width, height, and sequence length set to 512, 512, and 14, respectively. The denoising U-Net receives motion magnitude and frame rate, fixed at 127 and 8 respectively, as conditioning inputs. We initialize VAE and U-Net network parameters to pre-trained I2V-SVD (StabilityAI, 2023) parameters. During training, we freeze the VAE and fine-tune all U-Net parameters for the first 5K iterations. Afterward, we fix the temporal attention and temporal residual block layers in the U-Net and fine-tune for an additional 5K iterations.
Appendix B Sampling Strategy for TAP
@@
In this section, we show the effectiveness of our sampling strategy for TAP. Within the TAP scheme, we reduce computational cost by selecting 4 frames from the previous spatial tile, which consists of 14 frames. A straightforward approach is to select frames randomly. However, we opt for a more informed strategy: selecting frames whose keys exhibit large standard deviations. This choice is based on the observation that frames with more varied and sharper details tend to produce distinct keys, resulting in larger standard deviations. Figure 10 illustrates this trend by comparing random sampling with maximum standard deviation (max SD) sampling. As shown, max SD sampling produces consistent and detailed brick patterns, in contrast to random sampling. This demonstrates that selecting detail-rich frames based on their standard deviations enhances the quality and consistency of temporal tiles.
Appendix C Ablation Study of DSSAG
@@
Figure 11 demonstrates the effectiveness of DSSAG by comparing VSR results with and without DSSAG. Without DSSAG, the car grille and the bases of the windows exhibit significant distortions. However, after applying DSSAG, these elements are restored with clear and precise details. These results highlight the effectiveness of DSSAG in enhancing visual quality and structural fidelity in VSR.
Appendix D Image Generation Comparison of DSSAG
@@
Figure 10. Comparison of two different TAP sampling strategy.
Figure 11. Effects of DSSAG. The first row and second row are the results of DC-VSR without and with DSSAG. The proposed method enhances the generation quality of video frames.
@@
Table 3. Comparisons with previous guidance methods using Inception Score (IS) (Salimans et al., 2016) and Fréchet Inception Distance (FID) (Heusel et al., 2018) metrics. feed-forward/iter. indicates the number of feed-forward passes required by the diffusion model for each denoising iteration. The best and second-best scores are written as bold and underline.
@@
Appendix E Image-to-Video Stable Video Diffusion
where $D_{\theta}$ is a denoiser network, $\bm{c}$ is additional conditions such as text prompt, motion magnitude, and frame rate, $\sigma$ is sampled noise level from predefined $p_{train}(\sigma)$, $\bm{y}\sim p_{\mathcal{E}(data)}$, $\bm{n}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})$ and $\lambda(\sigma)$ is adaptive loss weight. For stable and consistent input and output signal magnitudes, $D_{\theta}$ is derived from network $F_{\theta}$ in the following form:
Appendix F tLP and tOF
Figure 12. Visualization of previous diffusion guidance approaches and our DSSAG. DSSAG combines well with CFG, while 1.5$\times$ faster than SAG (2023) and PAG (2024)

### v2

@@ 4.2. Component Evaluation
Effect of $\rho$ in DSSAG
DSSAG adjusts the detail suppression strength using $\rho$. Fig. 10 shows the effect of $\rho$. A smaller $\rho$ increases the blurring effect in Eq. 7, resulting in more enhanced edges. In contrast, a larger $\rho$ makes Eq. 7 closer to the conventional self-attention operation, producing less enhanced but more accurate details. This flexibility allows our framework to adapt to various scenarios: a small $\rho$ is suitable for severely degraded LR videos or animation videos, while a large $\rho$ works better for mildly degraded LR videos.
Table 2. Ablation study of SAP & TAP and comparison with previous guidance methods using the REDS4 dataset. $\text{DOVER}^{*}$ denotes DOVER ($\times$100). The best and second-best scores are marked in bold and underline. Our proposed DSSAG is 1.5×1.5$\times$1.5 × faster than SAG and PAG.
@@
Figure 6. Qualitative comparison with previous guidance approaches. The input video is from the VideoLQ dataset (Chan et al., 2022b) (sample 041).
@@ 5. Conclusion
Since DC-VSR is based on a diffusion model, it may generate false details and face difficulties in achieving real-time VSR. Moreover, DC-VSR may struggle to maintain temporal consistency across distant temporal tiles in long video clips, as TAP primarily relies on nearby tiles. Addressing these issues could be a promising future direction.
Acknowledgements.
This work was supported by Samsung Electronics Co., Ltd., the Institute of Information & Communications Technology Planning & Evaluation (IITP) grants (RS-2019-II191906, Artificial Intelligence Graduate School Program (POSTECH), No.2021-0-02068, Artificial Intelligence Innovation Hub), and the National Research Foundation of Korea (NRF) grant (No.2018R1A5A1060031), funded by the Korea government (MSIT).
@@ Acknowledgements.
We uploaded a video as Supplementary Material, containing VSR results and comparisons that include the contents mentioned in the main paper.
@@ Appendix A Implementation Details
We fine-tune Image-to-Video Stable Video Diffusion (I2V-SVD) (StabilityAI, 2023) with an 8×8$\times$8 × downsampling VAE (Rombach et al., 2022), trained on 14-frame sequences at a resolution of $576\times 1024$. We train DC-VSR using the AdamW optimizer (Loshchilov and Hutter, 2019) with a learning rate of $1\times 10^{-4}$ and a batch size of 256 on four NVIDIA A100-80GB GPUs. Due to the memory limit, we use gradient accumulation to match the size of the batch. The training data are cropped into spatio-temporal tiles with width, height, and sequence length set to 512, 512, and 14, respectively. The denoising U-Net receives motion magnitude and a frame rate, fixed at 127 and 8, respectively, as conditioning inputs. Additionally, the denoising U-Net takes a CLIP (Radford et al., 2021) embedding vector of the first frame of the low-resolution input video as a text prompt. We initialize VAE and U-Net network parameters to pre-trained I2V-SVD (StabilityAI, 2023) parameters. During training, we freeze the VAE and fine-tune all U-Net parameters for the first 5K iterations. Afterward, we fix the temporal attention and temporal residual block layers in the U-Net and fine-tune the U-Net for additional 5K iterations.
In our framework, Spatial Attention Propagation (SAP), Temporal Attention Propagation (TAP) and Detail-Suppression Self-Attention Guidance (DSSAG) are adopted to every spatial self-attention layer in the first two down-blocks and the last two up-blocks in the denoising U-Net. Merged keys and values in SAP and TAP are computed by concatenating each key and value along a flattened spatial dimension in the self-attention mechanism. In our experimental settings, DC-VSR is capable of performing video super-resolution (×$\times 4$) on input videos of up to 768$\times$768 resolution with 21 frames in a single NVIDIA A100-80GB GPU.
Appendix B Alternating and Simultaneous Strategies for SAP and TAP
While SAP and TAP may be applied simultaneously, our approach adopts an alternating application strategy to improve computational efficiency. Table 3 provides a performance comparison of the two approaches on the VideoLQ (Chan et al., 2022b) dataset. Both approaches show similar performance in MUSIQ (Xie et al., 2022) and $\text{DOVER}^{*}$ (Wu et al., 2023), but the simultaneous approach requires twice the inference time of the alternating approach. Therefore, we adopt the alternate approach in our proposed DC-VSR to ensure computational efficiency.
Figure 11. Comparison of two different TAP sampling strategy. The input video is from the VideoLQ dataset (Chan et al., 2022b) (sample 013).
Table 3. Comparison of alternating and simultaneous application of SAP and TAP on the VideoLQ (Chan et al., 2022b) dataset. The best and second-best scores are marked in bold and underline.
  Approach Alternating Simultaneous
MUSIQ $\uparrow$ 58.15 58.14
$\text{DOVER}^{*}$uparrow 78.30 78.31
Appendix C Sampling Strategy for TAP
@@
In this section, we show the effectiveness of our sampling strategy for TAP. Within the TAP scheme, we reduce the computational cost by selecting 4 frames from the previous spatial tile, which consists of 14 frames. A straightforward approach is to select frames randomly. However, we opt for a more informed strategy: selecting frames whose keys exhibit large standard deviations. This choice is based on the observation that frames with more varied and sharper details tend to produce distinct keys, resulting in larger standard deviations. Fig. 11 illustrates this trend by comparing random sampling with maximum standard deviation (max SD) sampling. As shown, max SD sampling produces consistent and detailed brick patterns, in contrast to random sampling. This demonstrates that selecting detail-rich frames based on their standard deviations enhances the quality and consistency of temporal tiles.
Appendix D Tile Propagation Range of TAP
In TAP, temporal information is propagated bidirectionally from both the previous and next tiles. While increasing the propagation range can further enhance temporal consistency, it also introduces higher computational overhead and raises the risk of propagating inconsistent information. Table 4 compares the impact of different propagation ranges in terms of warping error (WE). The results indicate that increasing the propagation range does not lead to significant performance improvements, as the directly adjacent tiles already provide sufficient information for TAP. Furthermore, although a propagation range of 2 achieves the best performance, further extending the range degrades performance due to the propagation of inconsistent information between distant tiles. In our experiments, we adopt a propagation range of 1, as the performance improvement from additional tiles is minimal compared to the substantial increase in memory usage, requiring 3.2GB of GPU memory per tile.
Table 4. Warping error (WE) as a function of the tile propagation range of TAP on the VideoLQ (Chan et al., 2022b) datasets (no.40-49). The best and second-best scores are marked in bold and underline, respectively.
  # Tiles 1 2 3 4
WE $\downarrow$ 0.1749 0.1744 0.1745 0.1746
Appendix E Subsampling Rates of SAP and TAP
SAP and TAP leverage subsampled information from the current frame and previous temporal tile to improve spatial and temporal consistency. A low subsampling rate, which preserves more information, can lead to improved performance, but significant computational costs are required. Table 5 shows correlation between subsampling rates and performance changes in SAP. We compare the average similarity over 28 frames between two distant spatial tiles, as indicated in Fig. 5 of our main paper. As the subsampling rate decreases, LPIPS (Zhang et al., 2018a), ST-LPIPS (Ghildyal and Liu, 2022), and CLIP similarity (Radford et al., 2021) scores improve, indicating the performance increase. Consequently, utilizing the full information from the current frame (100% in Table 5) yields the best results.
Table 6 shows warping error (WE) under varying subsampling rates of TAP on the VideoLQ (Chan et al., 2022b) dataset (no.40-49). Similar to the trend in SAP, sampling more key-value pairs, which correspond to more extensive use of previous tile information, enhance temporal consistency. In our experimental settings, we select the subsampling rate of 2 for SAP, and 4 key-value pairs for TAP, which correspond to $50\%$ and $28\%$ sampling, respectively, considering computational overhead and performance improvement.
Appendix F Ablation Study of DSSAG
@@
Fig. 12 demonstrates the effectiveness of DSSAG by comparing VSR results with and without DSSAG. Without DSSAG in Fig. 12, the car grille and the bases of the windows exhibit significant distortions. However, after applying DSSAG, these elements are restored with clear and precise details. These results highlight the effectiveness of DSSAG in enhancing visual quality and structural fidelity in VSR.
Table 5. Performance comparison based on SAP subsampling rates. The best and second-best scores are marked in bold and underline, respectively.
  Subsampling rate 8 (12.5%) 4 (25%) 2 (50%) 1 (100%)
LPIPS $\downarrow$ 0.4693 0.4551 0.4416 0.4303
ST-LPIPS $\downarrow$ 0.3065 0.2846 0.2641 0.2598
CLIP Similarity $\uparrow$ 0.9500 0.9574 0.9618 0.9637
Table 6. Performance comparison based on TAP subsampling rates. The top row represents the number of subsampled key-value pairs out of 14 key-value pairs from the previous temporal tile. The best and second-best scores are marked in bold and underline, respectively.
  # key-value pairs 1 2 4 8
WE $\downarrow$ 0.1767 0.1758 0.1749 0.1742
Appendix G Image Generation Comparison of DSSAG
@@
Appendix H Inference Time Comparison
Table 8 presents a computation time comparison of diffusion-based video super-resolution methods for 4×4$\times$4 × super-resolution of $640\times 360$ videos. Although DC-VSR significantly outperforms the previous approaches, it incurs slightly higher computational overhead, mainly due to the need for spatio-temporal tile splitting and merging at each diffusion timestep and attention-based approaches for maintaining spatial and temporal consistency.
Appendix I Image-to-Video Stable Video Diffusion
where $D_{\theta}$ is a denoiser network, $\bm{c}$ is additional conditions such as text prompt, motion magnitude, and frame rate, $\sigma$ is sampled noise level from predefined $p_{train}(\sigma)$, $\bm{y}\sim p_{\mathcal{E}(data)}$, $\bm{n}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})$ and $\lambda(\sigma)$ is adaptive loss weight. Here, CLIP (Radford et al., 2021) embedding vectors of the input image is given as text prompt. For stable and consistent input and output signal magnitudes, $D_{\theta}$ is derived from network $F_{\theta}$ in the following form:
Figure 12. Effects of DSSAG. The first row and second row are the results of DC-VSR without and with DSSAG. The proposed method enhances the generation quality of video frames. The input videos are from the VideoLQ dataset (Chan et al., 2022b) (sample 024, 039, and, 042).
@@ Appendix I Image-to-Video Stable Video Diffusion
Appendix J tLP and tOF
Table 7. Comparisons with previous guidance methods using Inception Score (IS) (Salimans et al., 2016) and Fréchet Inception Distance (FID) (Heusel et al., 2018) metrics. feed-forward/iter. indicates the number of feed-forward passes required by the diffusion model for each denoising iteration. The best and second-best scores are marked inbold and underline.
@@
Table 8. Diffusion inference time (in seconds/frame) comparison with previous diffusion-based video super-resolution tasks on 4×4$\times$4 × upscaling 640$\times$360 video.
  Methods UAVideo (2024) MGLD (2024a) Ours
Time (sec./frame) 31 54 67
Figure 13. Visualization of previous diffusion guidance approaches and our DSSAG. DSSAG combines well with CFG, while 1.5$\times$ faster than SAG (2023) and PAG (2024). Images are generated using a diffusion model (CompVis, 2022).